OpenStudy (anonymous):

Can some one help me solve this? ;) 3x – 9y = 3 6x – 3y = -24 i have no clue what to do. :(

OpenStudy (whpalmer4):

Two basic routes, substitution or elimination: Substitution is good if one of the equations is easy to arrange in terms of the other. For example, you have \[3x-9y=3\] which we solve for x and get \[3x=9y+3\]\[x=3y+1\]. Now we substitute 3y+1 wherever we see x in the other equation: \[6(3y+1)-3y=-24\]\[18y+6-3y=-24\]\[15y=-30\]\[y=-2\] Put y=-2 in either of the original equations and solve for x. \[3x-9(-2)=3\]\[3x+18=3\]\[3x=3-18=-15\]\[x=-5\] Important: put your answers back in one of the original equations and make sure they work! 3(-5)-9(-2) = -15+18 = 3, just like it should.

OpenStudy (whpalmer4):

Elimination works best if you can multiply one of the equations by some number that will make the coefficients be equal but opposite. If we multiply your first equation by -2 on both sides, we get \[-6x + 18y = -6\] and if we add that to the second equation, the -6x in the first equation and the 6x in the second equation cancel each other out and we are left with just y and numbers, which we can solve for y: \[-6x + 18y + 6x -3y = -6 - 24\] (I just added the left hand sides of the equations together and the right hand sides together) \[15y = -30\] \[y=-2] Then we plug that into one of the equations just like we did in substitution to find x.

OpenStudy (whpalmer4):

Oops, I made a typo, that should be \[15y = -30\]\[y=-2\]

OpenStudy (whpalmer4):

Now you should have some clues about how to do these :-)

OpenStudy (anonymous):

Thank you so so so so so much! You helped me so much! :D

OpenStudy (whpalmer4):

They are pretty simple to do, just takes some practice. Be sure to pay attention when multiplying negative numbers or subtracting expressions containing - signs — those are easy places to make mistakes.

OpenStudy (anonymous):

oh okay! thank you again! :)